A ball is thrown vertically upward. It has a speed of 10 m/sec when it has reached one half of its maximum height. How high does the ball rise?


Take g = 10 m/s2


(a)    10 m        


(b)    5 m


(c)    15m


             (d)    20m

Grade:10

2 Answers

Jitender Pal
10 years ago

(a)
For part AB
From 3rd equation of motion v2 = u2 - 2gH
 
0=u^2-2g (H/2)=u^2-gH
H=   u^2/g=  ?10?^2/10   =10 m

ankit singh
3 years ago

For part AB
From 3rd equation of motion v2 = u2 - 2gH
 
0=u^2-2g (H/2)=u^2-gH
H=   u^2/g=  ?10?^2/10   =10 m